# Effective bias expansion for circumventing 21 cm foregrounds

### v1

@@ Acknowledgments
It is a pleasure to thank Simon Foreman and Julian Muñoz for helpful discussions and comments on this paper. W.Q. was supported by the National Science Foundation Graduate Research Fellowship under Grant No. 2141064 and a grant from the Simons Foundation (Grant Number SFI-MPS-SFJ-00006250, W.Q.). K.-F.C. acknowledges support from the Mitacs Globalink Research Award, Taiwan Think Global Education Trust Scholarship, and the Taiwan Ministry of Education’s Government Scholarship to Study Abroad. K.S. acknowledges support from a Natural Sciences and Engineering Research Council of Canada (NSERC) Subatomic Physics Discovery Grant and from the Canada Research Chairs program. A.L. acknowledges support from an NSERC Discovery Grant and a Discovery Launch Supplement, and the William Dawson Scholarship at McGill. This analysis made use of Numpy Harris et al. (2020), Scipy Virtanen et al. (2020), Jupyter Kluyver et al. (2016), tqdm da Costa-Luis (2019), and Matplotlib Hunter (2007).

### v2

@@ Acknowledgments
It is a pleasure to thank Simon Foreman and Julian Muñoz for helpful discussions and comments on this paper. W.Q. was supported by the National Science Foundation Graduate Research Fellowship under Grant No. 2141064 and a grant from the Simons Foundation (Grant Number SFI-MPS-SFJ-00006250, W.Q.). K.-F.C. acknowledges support from the Mitacs Globalink Research Award, Taiwan Think Global Education Trust Scholarship, and the Taiwan Ministry of Education’s Government Scholarship to Study Abroad. K.S. acknowledges support from a Natural Sciences and Engineering Research Council of Canada (NSERC) Subatomic Physics Discovery Grant and from the Canada Research Chairs program. A.L. acknowledges support from an NSERC Discovery Grant and a Discovery Launch Supplement, and the William Dawson Scholarship at McGill. This analysis made use of Numpy Harris and others (2020), Scipy Virtanen et al. (2020), Jupyter Kluyver and others (2016), tqdm da Costa-Luis (2019), and Matplotlib Hunter (2007).